Find f + g, f − g, f * g, and f/g.
Also specify the domain for each.
f(x) = −5x − 9, g(x) = −3x + 10

Answer by CubeyThePenguin(3113)

(f+g)(x) = (-5x-9) + (-3x+10) = -8x+1

(f-g)(x) = (-5x-9) - (-3x+10) = -2x-19

(f*g)(x) = (-5x-9)(-3x+10) = 15x^2-23x-90

(f/g)(x) = (-5x-9)/(-3x+10)